OpenStudy (anonymous):

Convert 1 7/20 to a decimal?

OpenStudy (stefrheart):

is it like this?\[1 \frac{ 7 }{ 20 }\]

OpenStudy (anonymous):

1.35

OpenStudy (anonymous):

it shows as 1 7/20 =

OpenStudy (anonymous):

thank you.

OpenStudy (stefrheart):

do 27/20 and divide that on your calculator

OpenStudy (anonymous):

I did it with a calculator, but you know it has a whole 1 because of the initial 1 and \[\frac{ 7 }{ 20 } = .35\]